Marlin24k:

12x+9x+24=180 help me find x.

sxdsouls:

you wqnana start by simplifying both sides of the equation 12x+9x+24=180 comebine like terms with 24 (12x+9x)+(24)=180 Subtract 24 from both sides. 21x+24−24=180−24 divide both sides by 21 21x/21 = 156/21 \[x=52/7\]